5kg mass, moving at 3( m)/(s) has a force of 20 N applied in the same direction in which it is moving for 3 seconds, what will be the final velocity
1
Expert's answer
2019-02-01T11:07:08-0500

Since force F and mass m are given, the acceleration a can be calculated directly from Newton’s second law

F=ma

The magnitude of the acceleration is

a=F/m (1)

The magnitude of the velocity at a time t is determined by the kinematics equation

v=v_0+at, (2)

where v_0=3 m/s is the initial velocity. Substituting (1) into (2), yield

v=v_0+F/m t.

Plugging in known values, we have

v=3 m/s+(20 N)/(5 kg)⋅3 s=3 m/s+12 m/s=15 m/s